A rigid windshield wiper on a road repair vechile has a length of 14 inches. The rubber wiper itself is 12 inches long. How great an area does it wipe clean if the arc through which it moves measures 100 degrees?
Answer by fractalier(2101)

The area wiped clean would be 100/360 of the area of a full circle.
That would be
(100/360)(pi)(r^2) =
(100/360)(pi)(12^2) =
(14400/360)(pi) =
40(pi) sq in